OpenStudy (anonymous):

At the movie theatre, child admission is $5.30 and adult admission is $9.00 On Tuesday, three times as many adult tickets as child tickets were sold, for a total sales of $710.60 How many child tickets were sold that day?

OpenStudy (anonymous):

@bhaskar.pathk HELP PLEASE lol i need this ive been doing this hw for hours i still have others left

OpenStudy (anonymous):

@jocelynn : take it simple way... let the number of adult ticket sold be a and number of child ticket be c then as per qstn 53.30 x c+ 9.00 x a = 710.60

OpenStudy (anonymous):

and since "three times as many adult tickets as child tickets were sold" so a=3c now solve the equation..

OpenStudy (anonymous):

can you draw it so i can see the way your trying to put it? please @bhaskar.pathk

OpenStudy (anonymous):

\[53.3 c+9 time 3c=710.60\] \[53.3c+27c=710.60 \] 80.3c=710.60

OpenStudy (anonymous):

can you recheck fhe data in the question or post the link

OpenStudy (anonymous):

yes, the data in the question is correct @bhaskar.pathk

OpenStudy (anonymous):

the aswer is coming out as a decimal ... i dont believe its suppose to be a decimal ? @bhaskar.pathk

OpenStudy (anonymous):

@jocelynn : that's why said to rechek the data... any way let me retry..

OpenStudy (anonymous):

thank you sooo much, your a life saver @bhaskar.pathk